# Bounds on the number of squares in recurrence sequences

### v1

@@ 3.2. Lower bounds for $y_{k}$’s
(b) For all $k$, $2y_{k}$ is a positive integer. The sequences $\left(y_{k}\right)_{k\geq 0}$ and $\left(y_{k}\right)_{k\leq K+1}$ are increasing sequences of positive numbers.
@@
(3.13) $y_{k}>\left\{\begin{array}[]{ll}\left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{k-1}&\text{for $k>0$,}\\ \left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{K-k}&\text{for $k\leq K$.}\end{array}\right.$
@@
If $N_{\alpha}>0$, then $y_{k}<0$ can occur for $k<0$.
@@
Writing $\varepsilon^{2k}=t_{k}+u_{k}\sqrt{d}$ for $k\geq 1$, we see that $\left(u_{k}\right)_{k\geq 1}$ is an increasing sequence. Hence
$4b^{2}y_{|K|}$ $=b^{4}\left(t_{|K|}^{2}+du_{|K|}^{2}\right)-2ab^{2}t_{|K|}u_{|K|}=\left(b^{2}t_{|K|}-au_{|K|}\right)^{2}-N_{\alpha}u_{|K|}^{2}$
@@
So equation (3.13) holds for $k=K$.
Recall the recurrence in (1.3) for $k\geq 0$ and using the monotonicity established in part (b):
@@
If $N_{\varepsilon}=1$, then the stronger form of the desired inequality holds, so we need only consider $N_{\varepsilon}=-1$. We always have $du^{2}\geq 5$ (with equality only if $(d,t,u)=(5,1,1)$. Otherwise, $du^{2}\geq 8$). Hence the desired inequality holds.
This is not too far from the best possible result from this argument. For $(a,d,t,u)=(1,5,1,1)$ and $b$ arbitrarily large, $y_{2}/y_{1}$ approaches $7/3$ from above, whereas we use $2$ here.
Using the recurrence $y_{k-1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k+1}$, we obtain the desired result for $k\leq K$ too. ∎
@@ 3.3. Gap Principle
Multiplying by $\omega^{-1/4}$ and expanding the resulting expression, the product above equals
@@
$=\frac{2-\cos(4\varphi)}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}.$
Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $016\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=$ 0 and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$. Part (b) now follows and with inequalities since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$. ∎
@@
(a) Suppose that $-N_{\alpha}$ is a square. If $y_{i}$ and $y_{j}$ are both squares with $y_{j}>y_{i}\geq 4b^{6}$, then
$y_{j}>\frac{57.6d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
(b) Suppose $-N_{\alpha}$ is not a square. If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4b^{6}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
@@
If $y_{i}\geq 4.3b^{6}\left|N_{\alpha}\right|$, then we can replace $15.36$ with $182$.
@@
By Proposition 3.1, there are integers $f_{j}$, $r_{j}$ and $s_{j}$, and choices of signs such that
@@
where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$.
For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $y_{i},y_{j}>1$, we will assume that the $\pm$ on the left-hand side is always $+$. That is,
@@
$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
@@
$=\frac{2b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}\,y_{j}}<\frac{2b^{2}}{y_{j}}\leq\frac{1}{2},$
the last inequality holding because $y_{j}\geq 4b^{6}$.
@@
From (3.19), we can apply Lemma 3.6(b) with $c_{0}=1/2$ to obtain
@@
In the same way, we define $\zeta_{4}^{(i)}$ for any square $y_{i}>1$ in our sequence and (3.21) also holds with $j$ replaced by $i$ for such $i$. Hence, for any two distinct squares, $y_{i},y_{j}\geq 4b^{2}$, among the $y_{k}$’s, we have
@@
(3.24) $\begin{array}[]{rl}-2\zeta_{4}^{(i)}y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}&\text{if $\zeta_{4}^{(j)}=\zeta_{4}^{(i)}$,}\\ 2\zeta_{4}^{(i)}x&\text{if $\zeta_{4}^{(j)}=-\zeta_{4}^{(i)}$,}\\ \zeta_{4}^{(i)}\left(1-\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=\sqrt{-1}\zeta_{4}^{(i)}$ and}\\ \zeta_{4}^{(i)}\left(1+\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=-\sqrt{-1}\zeta_{4}^{(i)}$.}\end{array}$
@@
(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that $1+\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
@@
(3.27) $y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3},$
We can use this gap principle to improve its constant term. Combining (3.27) with $y_{i}\geq 4b^{6}>4b^{2}\left|N_{\alpha}\right|/d$ and $\left|N_{\alpha}\right|<b^{4}d$, we obtain
$y_{j}>\frac{3.84}{b^{12}}y_{i}^{3}\geq 61.44y_{i},$
@@
$\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\frac{1+1/61.44}{y_{i}}.$
@@
$y_{j}>\frac{57.61}{b^{12}}\left(\frac{a^{2}+\left|N_{\alpha}\right|}{\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
completing our proof of part (a).
@@
Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is ±$\pm 1$ times the other. Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $i<j$.
@@
$y_{j}>15.36\left(\frac{b^{2}d}{f_{i}f_{j}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
@@
Since $f_{i}f_{j}\leq\left(4b^{2}\right)^{2}\left|N_{\alpha}\right|\leq 16b^{8}d$, if $y_{i}\geq 4.3b^{6}\left|N_{\alpha}\right|$, then $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, as required. ∎
@@ 3.4. Miscellaneous Lemmas
(a) Let $y_{k}\geq 4b^{2}$ be a square and put
@@
(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4b^{2}$, $\omega$ be as in part (a) and put $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$ with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1. Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.7 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$. Then
@@
$x_{k}^{2}+N_{\alpha}$ $=dy_{k}^{2}+2N_{\alpha}>d\left(y_{k}^{2}-2b^{4}\right)\geq dy_{k}^{2}\left(1-\frac{2b^{4}}{y_{k}^{2}}\right)$
$\geq 0.875dy_{k}^{2}>0.875x_{k}^{2},$
since $y_{k}\geq 4b^{2}$ and hence $1-2b^{4}/y_{k}^{2}\geq 1-2b^{4}/\left(16b^{12}\right)\geq 0.875$.
@@
Since $y_{k}\geq 4b^{2}$, we have $dy_{k}^{2}\geq 16db^{4}>16\left|N_{\alpha}\right|$ and so
$x_{k}^{2}=dy_{k}^{2}+N_{\alpha}>15\left|N_{\alpha}\right|,$
which yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
@@
$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{N_{\alpha}}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{i}-s_{i}\sqrt{N_{\alpha}}\right)^{4}=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
@@
$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$\frac{\left(r_{k}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{k}^{2}\right)^{4}\left(\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right)}{f_{k}^{2}}$
$f_{k}^{2}y_{k}^{2}\left(\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right)$.
Applying equation (3.3) with $j=\ell$ and with $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ as defined in the statement of this lemma, we have
(3.29) $|f(x,y)|$ $=\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
$=2f_{k}^{2}f_{\ell}^{2}y_{k}^{2}\sqrt{\left|N_{\alpha}\right|}.$
@@
(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$ $=\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
$=\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}<\frac{2b^{2}}{y_{\ell}}\leq\frac{1}{2}.$
since $y_{\ell}\geq 4b^{2}$.
By Lemma 3.6(b) with $c_{0}=1/2$,

### v2

@@ 3.2. Lower bounds for $y_{k}$’s
(b) For all $k$, $2y_{k}$ is a positive integer. The sequences $\left(y_{k}\right)_{k\geq 0}$ and $\left(y_{K+1},y_{K},y_{K-1},y_{K-2},\ldots\right)$ are increasing sequences of positive numbers.
@@
(3.13) $y_{k}\geq\left\{\begin{array}[]{ll}\left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{k-1}&\text{for $k>0$,}\\ \left(\left|N_{\alpha}\right|u^{2}/\left(4b^{2}\right)\right)\left(2du^{2}/5\right)^{\max(0,K-k)}&\text{for $k<0$.}\end{array}\right.$
@@
The condition $N_{\alpha}<0$ is needed, since if $N_{\alpha}>0$, then $y_{k}<0$ can occur for $k<0$.
Also, since $du^{2}\geq 5$ holds under the conditions here, the lower bounds in (3.13) are increasing as $|k|$ increases.
@@
Writing $\varepsilon^{\ell}=t_{\ell}+u_{\ell}\sqrt{d}$ for $\ell\geq 1$ (note that $u_{1}=u$ in the notation of this lemma), we can show by an easy induction that $\left(u_{\ell}\right)_{\ell\geq 1}$ is an increasing sequence. Hence, for $k<0$,
$4b^{2}y_{k}$ $=b^{4}\left(t_{|2k|}^{2}+du_{|2k|}^{2}\right)-2ab^{2}t_{|2k|}u_{|2k|}=\left(b^{2}t_{|2k|}-au_{|2k|}\right)^{2}-N_{\alpha}u_{|2k|}^{2}$
@@
So equation (3.13) holds for $k$ satisfying $K\leq k<0$.
Recalling the recurrence in (1.3) for $k\geq 0$ and using the monotonicity established in part (b), we obtain
@@
If $N_{\varepsilon}=1$, then the stronger form of the desired inequality stated after equation (3.13) in the lemma holds, so we need only consider $N_{\varepsilon}=-1$. We always have $du^{2}\geq 5$ (with equality only if $(d,t,u)=(5,1,1)$. Otherwise, $du^{2}\geq 8$, which we use to establish the other inequality after equation (3.13) in the lemma). Hence the desired inequality holds.
Using the recurrence $y_{k-1}=\left(du^{2}+2N_{\varepsilon}\right)y_{k}-y_{k+1}$, we obtain equation (3.13) for $k\leq K$ in the same way. ∎
We will also need to know when $K$ in Lemma 3.5 is not equal to $-1$ for $b=1$.
Lemma 3.6.
[…]
We have $y_{-1}>1$, except for
(i) $a\geq 1$, $d=a^{2}+4$, $t=a$ and $u=1$, where $\alpha=2\varepsilon$ and $N_{\alpha}=-4$,
(ii) $a\geq 1$, $d=a^{2}+1$, $t=2a$ and $u=2$, where $\alpha=\varepsilon$ and $N_{\alpha}=-1$.
In these cases, we have $y_{-2}=a^{2}+1$ and $y_{-2}=4a^{2}+1$, respectively. So in each of these cases, $K=-2$.
Proof.
From (3.13), we have $y_{1}\geq\left|N_{\alpha}\right|u^{2}/4$. So if $y_{-1}=1$, we must have either (1) $N_{\alpha}=-4$ and $u=1$, or (2) $N_{\alpha}=-1$ and $u=2$.
In case (1), we have $d=t^{2}\pm 4$ and $a^{2}-d=-4$. Substituting the expression for $d$ into the second equation, we get $a^{2}-t^{2}=-4\pm 4=0,-8$.
$a^{2}-t^{2}=-8$ can only occur for $a=1$ and $t=3$ (since we are assuming that $a$ and $t$ are positive integers). But in this case, $b^{2}t-au=2$, so $84y_{-1}=\left(b^{2}t-au\right)^{2}-N_{\alpha}u^{2}=$ 8.
Otherwise, we have $t=a$. So $d=a^{2}+4$, $\alpha=2\varepsilon$ and $N_{\varepsilon}=-1$. Here $b^{2}t-au=0$, so $y_{-1}=1$.
In case (2), we have $d=\left(t^{2}\pm 4\right)/4$ and $a^{2}-d=-1$. We proceed similarly, obtaining $84a^{2}-t^{2}=-$ 8 (which is never possible) when $N_{\varepsilon}=1$ and $04a^{2}-t^{2}=$ 0 when $N_{\varepsilon}=-1$. Here we have $t=2a$ and $d=a^{2}+1$, so $\alpha=\varepsilon$. Once again, $b^{2}t-au=0$, so $y_{-1}=1$.
A direct calculation, done using Maple, provides the values of $y_{-2}$. ∎
@@ 3.3. Gap Principle
Multiplying by $\omega^{-3/4}$ and expanding the resulting expression, the product above equals
@@
$=\frac{2-2\cos(4\varphi)}{8\cos^{2}(\varphi)(\cos(\varphi)+1)}.$
Since $\cos(4\varphi)=8\cos^{4}(\varphi)-8\cos^{2}(\varphi)+1$, we have $016\cos^{4}(\varphi)-16\cos^{2}(\varphi)+c_{0}^{2}=$ 0 and using this relation, we find that $c_{0}^{2}/\left(8\cos^{2}(\varphi)(\cos(\varphi)+1)\right)$ is a root of the polynomial $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$. Part (b) now follows and follows with inequalities too, since the smallest positive real root of $x^{4}-8x^{3}+20x^{2}-16x+c_{0}^{2}$ grows with $c_{0}$. ∎
@@
(a) Suppose that $-N_{\alpha}$ is a square. If $y_{i}$ and $y_{j}$ are distinct squares with $i,j\neq 0$ and $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},b^{2}\left|N_{\alpha}\right|/d\right\}$, then
$y_{j}>57.32\left(\frac{d}{b^{2}\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}.$
(b) Suppose $-N_{\alpha}$ is not a square. If $y_{k_{1}}$, $y_{k_{2}}$ and $y_{k_{3}}$ are three distinct squares with $k_{1},k_{2},k_{3}\neq 0$ and $y_{k_{3}}>y_{k_{2}}>y_{k_{1}}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, then there exist distinct $i,j\in\{k_{1},k_{2},k_{3}\}$ such that
@@
If $y_{j}>y_{i}\geq\max\left\{4\sqrt{\left|N_{\alpha}\right|/d},4.27b^{2}\left|N_{\alpha}\right|^{2}/d\right\}$, then we can replace $15.36$ with $182$.
@@
Since we have assumed that $i\neq 0$ and $j\neq 0$, we can apply Proposition 3.1 to show that there are integers $f_{j}$, $r_{j}$ and $s_{j}$, and choices of signs such that
@@
where $f_{j}$ satisfies $f_{j}|\left(4b^{2}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)\right)$ and $f_{j}^{2}<16b^{4}\operatorname{sf}\left(\left|N_{\alpha}\right|\right)$. Recall that the other terms in the relationships for $f_{j}$ in Proposition 3.1 are not present here since $N_{\varepsilon}=\pm 1$.
For any two distinct squares among the $y_{k}$’s, say $y_{i}$ and $y_{j}$ with $i\neq 0$ and $j\neq 0$, we will assume that the $\pm$ on the left-hand side of (3.3) is always $+$. That is,
@@
(3.18) $\left(a+\sqrt{N_{\alpha}}\right)\left(r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
@@
$=\frac{2b^{2}\sqrt{\left|N_{\alpha}\right|}}{\sqrt{a^{2}+\left|N_{\alpha}\right|}\,y_{j}}\leq\frac{1}{2},$
the last inequality holds because $a^{2}+\left|N_{\alpha}\right|=db^{4}$ and $y_{j}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$.
@@
From (3.19), we can apply Lemma 3.7(b) with $c_{0}=1/2+0.0001$ to obtain
@@
In the same way, we define $\zeta_{4}^{(i)}$ for any square $y_{i}$ with $i\neq 0$ in our sequence and (3.21) also holds with $j$ replaced by $i$. Hence, for any two distinct squares, $y_{i},y_{j}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, among the $y_{k}$’s, we have
(3.22) $\left|\zeta_{4}^{(i)}\frac{r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}-\zeta_{4}^{(j)}\frac{r_{j}+s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{r_{j}-s_{j}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right|$
@@
(3.24) $\begin{array}[]{rl}-2\zeta_{4}^{(i)}y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}&\text{if $\zeta_{4}^{(j)}=\zeta_{4}^{(i)}$,}\\ 2\zeta_{4}^{(i)}x&\text{if $\zeta_{4}^{(j)}=-\zeta_{4}^{(i)}$,}\\ \zeta_{4}^{(i)}\left(1-\sqrt{-1}\right)\left(x-y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=\sqrt{-1}\zeta_{4}^{(i)}$ and}\\ \zeta_{4}^{(i)}\left(1+\sqrt{-1}\right)\left(x+y\sqrt{-\operatorname{sf}\left(N_{\alpha}\right)}\right)&\text{if $\zeta_{4}^{(j)}=-\sqrt{-1}\zeta_{4}^{(i)}$.}\end{array}$
@@
(a) Suppose that $-N_{\alpha}$ is a square. From (3.24), we see that at least one of $1\pm\sqrt{-1}$ always divides the numerator of the right-hand side of (3.23) and so
@@
(3.27) $y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
We can use this gap principle to improve its constant term. Combining (3.27) with $y_{i}\geq b^{2}\left|N_{\alpha}\right|/d$, we obtain
$y_{j}>\frac{3.84d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}\left(\frac{b^{2}\left|N_{\alpha}\right|}{d}\right)^{2}y_{i}=3.84y_{i}.$
@@
$\frac{\sqrt{2}}{b\left(y_{i}y_{j}\right)^{1/4}}<0.5051b^{2}\sqrt{\frac{\left|N_{\alpha}\right|}{a^{2}+\left|N_{\alpha}\right|}}\frac{1+1/3.84}{y_{i}}.$
@@
$y_{j}>\frac{24.34}{b^{12}}\left(\frac{a^{2}+\left|N_{\alpha}\right|}{\left|N_{\alpha}\right|}\right)^{2}y_{i}^{3}=\frac{24.34d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
Repeating the process from (3.27) onwards with $3.84$ replaced by $24.34$, we obtain
$y_{j}>\frac{52.31d^{2}}{b^{4}\left|N_{\alpha}\right|^{2}}y_{i}^{3}.$
Repeating it again with $52.31$ instead of $3.84$, we improve the constant to $56.97$. Repeating it one final time with $56.97$, we obtain the inequality in part (a).
@@
Among the three values $\zeta_{4}^{(k_{1})}$, $\zeta_{4}^{(k_{2})}$ and $\zeta_{4}^{(k_{3})}$, there must be at least one pair such that one member of the pair is ±$\pm 1$ times the other. Choose any such pair and let $i$ and $j$ be the associated indices, ordered so that $y_{i}<y_{j}$.
@@
which establishes the first lower bound for $y_{j}$ in part (b).
@@
Since $f_{i}f_{j}\leq\left(4b^{2}\right)^{2}\left|N_{\alpha}\right|$, if $y_{i}\geq 4.27b^{2}\left|N_{\alpha}\right|^{2}/d$, then $y_{i}>f_{i}f_{j}\left|N_{\alpha}\right|\sqrt{1.09/15.36}/\left(b^{2}d\right)$, as required. ∎
@@ 3.4. Miscellaneous Lemmas
[…]
(a) Let $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$ be a square and put
@@
(b) Let $y_{k}$ and $y_{\ell}$ be two squares with $y_{\ell}>y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, $\omega$ be as in part (a) and put
$x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}=\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)$
with $r_{k},r_{\ell},s_{k},s_{\ell}$ as in Proposition 3.1. Furthermore, suppose that the quantities $\zeta_{4}^{(k)}$ and $\zeta_{4}^{(\ell)}$ defined in (3.20) in the proof of Lemma 3.8 satisfy $\zeta_{4}^{(k)}=\pm\zeta_{4}^{(\ell)}$. Then
@@
$x_{k}^{2}+N_{\alpha}=dy_{k}^{2}+2N_{\alpha}=dy_{k}^{2}\left(1+\frac{2N_{\alpha}}{dy_{k}^{2}}\right)\geq 0.875dy_{k}^{2}>0.875x_{k}^{2},$
since $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$ and $x_{k}^{2}-dy_{k}^{2}=N_{\alpha}<0$.
@@
Since $y_{k}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$, we have $dy_{k}^{2}\geq 16\left|N_{\alpha}\right|$ and so
$x_{k}^{2}=dy_{k}^{2}+N_{\alpha}\geq 15\left|N_{\alpha}\right|.$
Combining this with the inequality above it yields $|\varphi|<2.29\sqrt{\left|N_{\alpha}\right|}/\left|x_{k}\right|<2.29/\sqrt{15}<0.6$.
@@
$\left(a+\sqrt{N_{\alpha}}\right)\left(r_{i}+s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{i}-s_{i}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
=±$=\pm 2f_{i}^{2}\sqrt{N_{\alpha}}$
@@
$\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}-s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(r_{k}+s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}$
$=$ $\frac{\left(r_{k}^{2}-\operatorname{sf}\left(N_{\alpha}\right)s_{k}^{2}\right)^{4}}{f_{k}^{2}}\left[\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
$\hskip 93.89409pt\left.-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right]$
$=$ $\frac{f_{k}^{2}y_{k}^{2}}{b^{4}}\left[\left(a+\sqrt{N_{\alpha}}\right)\left(r_{\ell}+s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
$\hskip 28.45274pt\left.-\left(a-\sqrt{N_{\alpha}}\right)\left(r_{\ell}-s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right].$
Applying equation (3.18) with $j=\ell$ to the last expression and with $x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}$ as defined in the statement of this lemma, we have
(3.29) $|f(x,y)|=$ $\left|\left(x_{k}+N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right.$
$\left.-\left(x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right)\left(x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right)^{4}\right|$
$=$ $\frac{2f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}}\sqrt{\left|N_{\alpha}\right|}.$
@@
(3.30) $\left|\omega-\left(\frac{x-y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}{x+y\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}}\right)^{4}\right|$
$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}f_{k}^{2}f_{\ell}^{2}y_{k}^{2}}{b^{4}\left|x_{k}-N_{\varepsilon}\sqrt{N_{\alpha}}\right|\left|r_{k}\mp s_{k}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}\left|r_{\ell}\mp s_{\ell}\sqrt{\operatorname{sf}\left(N_{\alpha}\right)}\right|^{4}}$
$=$ $\frac{2\sqrt{\left|N_{\alpha}\right|}}{\sqrt{d}\,y_{\ell}}\leq\frac{1}{2}.$
since $y_{\ell}\geq 4\sqrt{\left|N_{\alpha}\right|/d}$.
By Lemma 3.7(b) with $c_{0}=1/2+0.0001$,